OpenStudy (anonymous):

Calculus Help? Find any critical numbers of the function h(x)=sin^2(x)+cosx and the limits are 0

OpenStudy (anonymous):

I have no idea how to do this after finding the derivative so any help would be great!

OpenStudy (anonymous):

critical numbers include when h'(x) =0 and when h''(x) = 0

OpenStudy (anonymous):

Are regular zeros also critical numbers? I don't remember...

OpenStudy (anonymous):

What did you get for h' ?

OpenStudy (anonymous):

yeah I got that part down but how would I exactly use it, the derivative that I got was f'(x) = 2(sinx)(cosx)-(sinx)

OpenStudy (anonymous):

ok, set that equal to zero: 0 = 2(sinx)(cosx)-(sinx)

OpenStudy (anonymous):

factor out sin(x): 0 = 2(sinx)(cosx)-(sinx) 0 = sin(x)(2cosx - 1) and solve the two parentheses for zero: sinx = 0 and 2cosx - 1 = 0

OpenStudy (anonymous):

ahh so would I need to take the inverse of sin for sinx = 0 then?

OpenStudy (anonymous):

which would equal 0?

OpenStudy (anonymous):

Well, yes, but I prefer to look at the unit circle, sinx = 0 at only one place in your domain ...(unless you meant to include 0)...

OpenStudy (anonymous):

lol yeah I got it from the unit circle

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

wait would that be 0? on the unit circle the sin of 0 is 0..?

OpenStudy (anonymous):

yes, and also pi. Now solve 2cosx-1 = 0

OpenStudy (anonymous):

so that would be cosx =1/2 which would be pi/3

OpenStudy (anonymous):

would these values be my answer or would I need to plug them in still to the original equation?

OpenStudy (anonymous):

And one more answer for cosx = 1/2

OpenStudy (anonymous):

5pi/3

OpenStudy (anonymous):

nevermind answered my own question, thanks a lot for guiding me through the problem! :)

OpenStudy (anonymous):

Good, those are all the critical numbers in the domain for the first derivative.

OpenStudy (anonymous):

thanks I understand them a lot better now